OpenStudy (anonymous):

how do you solve for slope intercept form of 3x+2y=6?

OpenStudy (aroub):

there is 2 ways to do it but i like to do this way: y=mx+b 2y=-3x+6 y=-3/2x+3

OpenStudy (anonymous):

try to write eq. as y=mx +c where m is slope and c is intercept of line on y -axis.

OpenStudy (aroub):

yeah! same thing :)

OpenStudy (aroub):

and there is a way by using the point slope formula

OpenStudy (anonymous):

THANK YOU!!!! IM OUT!! SEE YA!!